# The Hurwitz existence problem and the prime-degree conjecture: A computational perspective

### v1

@@ 3 Eliminating False Positives
$=$ $\frac{\frac{(n-1)!}{n_{1}!\dotsm n_{m}!}\cdot\left(\sum_{\nu\vdash d_{j}}\chi^{\mu_{j}}(C_{\nu})\,|C_{\nu}|\right)^{3}}{\prod_{j=1}^{n}\dim S^{\mu_{j}}\cdot(d_{j}!)^{2}}$ (4)
@@ 3.1 Bounding the coefficients of $F_{d,3}$
$\frac{\frac{(n-1)!}{n_{1}!\dotsm n_{m}!}}{\prod_{j=1}^{n}\dim S^{\mu_{j}}\cdot(d_{j}!)^{2}}.$ (5)
@@
$\frac{(n-1)!}{n_{1}!\dotsm n_{m}!}\in\mathbb{Z}_{>0},$
@@
it follows that the numerator of (4) can be bounded by $(d!)^{4}$, while the denominator can be bounded by $(d!)^{3}$.
@@
we may take $(d!)^{3}$ as a common denominator. With this denominator, the numerator of each summand corresponding to a $\omega$ is bounded by $(d!)^{7}$. Therefore, if there are $p_{2}(d)$ secondary partitions $\omega\vdash\vdash d$, summing over the number of $\omega$ we obtain
@@
After using a common denominator as $(d!)^{3}$, every coefficient of the polynomial $F_{d,3}$ is a rational number whose numerator (in absolute value) is at most
$p_{2}(d)\,\cdot(d!)^{7}.$
@@ Acknowledgements
B.X. would like to express his sincere gratitude to Yufei Bai, Xiaoyang Chen, Tielin Dai, Dun Liang, Sicheng Lu, Zijin Peng, Qingyu Xu, Junyi Yang, Wenyan Yang and Xiaoya Zhai for many valuable and stimulating conversations during the course of this project. The authors are also grateful to Chao Wu at the Supercomputing Center of USTC for kindly answering their questions on the supercomputing system there.

### v2

@@ 3 Eliminating False Positives
$=$ $\frac{\frac{n!}{n_{1}!\dotsm n_{m}!}\cdot\left(\sum_{\nu\vdash d_{j}}\chi^{\mu_{j}}(C_{\nu})\,|C_{\nu}|\right)^{3}}{n\cdot\prod_{j=1}^{n}\dim S^{\mu_{j}}\cdot(d_{j}!)^{2}}$ (4)
@@ 3.1 Bounding the coefficients of $F_{d,3}$
$\frac{\frac{n!}{n_{1}!\dotsm n_{m}!}}{n\cdot\prod_{j=1}^{n}\dim S^{\mu_{j}}\cdot(d_{j}!)^{2}}.$ (5)
@@
$\frac{n!}{n_{1}!\dotsm n_{m}!}\in\mathbb{Z}_{>0},$
@@
it follows that the numerator of (4) can be bounded by $(d!)^{4}$, while the denominator can be bounded by $n\cdot(d!)^{3}$.
@@
we may take $(d!)^{4}$ as a common denominator. With this denominator, the numerator of each summand corresponding to a $\omega$ is bounded by $(d!)^{8}$. Therefore, if there are $p_{2}(d)$ secondary partitions $\omega\vdash\vdash d$, summing over the number of $\omega$ we obtain
@@
After using a common denominator as $(d!)^{4}$, every coefficient of the polynomial $F_{d,3}$ is a rational number whose numerator (in absolute value) is at most
$p_{2}(d)\,\cdot(d!)^{8}.$
@@ Acknowledgements
B.X. would like to express his sincere gratitude to Yufei Bai, Xiaoyang Chen, Tielin Dai, Dun Liang, Sicheng Lu, Xin Nie, Zijin Peng, Qingyu Xu, Junyi Yang, Wenyan Yang, Xiaoya Zhai and Youliang Zhong for many valuable and stimulating conversations during the course of this project. The authors are also grateful to Chao Wu at the Supercomputing Center of USTC for kindly answering their questions on the supercomputing system there.